# Privacy Preserving Federated Unsupervised Domain Adaptation with Application to Age Prediction from DNA Methylation Data

### v1

@@ 3 Discussion
In this study we present a framework called freda, for privacy-preserving unsupervised domain adaptation on high dimensional data in a distributed setting, addressing a significant limitation of the state-of-the-art method, wenda, which requires that both source and target domain data are held by a single entity. Our experimental results demonstrate that freda performs comparably to wenda and significantly outperforms the non-adaptive baseline model en-ls, even when the source domain data is distributed across multiple parties. Notably, for the cerebellum samples, which present a challenging distribution shift problem, freda effectively handles the shift, yielding comparable performance to that of wenda while ensuring complete data and model privacy in a distributed setting.
The significance of our findings lies in the ability of freda to maintain high predictive performance without compromising data and model privacy in a distributed environment. This is crucial for real-world applications, especially in fields like computational biology and medicine, where data is often sensitive and distributed across multiple institutions or hospitals. By leveraging federated learning and privacy-enhancing technologies, freda ensures that no private data is shared between parties while still addressing the distribution shift between the domains.
Our findings also contribute to the limited body of work on unsupervised domain adaptation for regression tasks in distributed scenarios, particularly in high-dimensional biological data. While unsupervised domain adaptation has been extensively studied for classification tasks, its application to regression, especially in privacy-sensitive contexts, remains less explored. The results from this article highlight the feasibility of adapting such techniques to real-world biological datasets and provide a strong baseline for further research in this area.
Our study has two limitations. First, although we demonstrate the effectiveness of freda on a publicly available and real-life dataset, the performance of our framework may vary with different types of data or in other application domains. Second, the reliance on available side information about tissue similarities, while not uncommon, may not always be practical in real-world scenarios where such information is limited or unavailable.
Future research should address the current limitations by exploring the application of freda to other types of data and domains, as well as investigating alternative methods for computing the optimal regularization parameter $\lambda$ without requiring prior knowledge about tissue similarities.
4 The Wenda Method
Wenda, proposed by Handl et al. [6], is the state-of-the-art solution which performs unsupervised domain adaptation on high dimensional tabular data. By extending traditional approaches that rely on the covariate shift assumption, the wenda method operates under the assumption that while the marginal distributions of inputs in the source domain $P_{S}(X)$ and target domain $P_{T}(X)$ may differ, the conditional distributions $P_{S}(Y|X)$ and $P_{T}(Y|X)$ are identical [6]. This classical assumption is known as the covariate shift assumption [23].
Wenda builds upon this framework by weakening the assumption to account for the possibility that some features might have different influences on the output in the source and target domains. Specifically, wenda assumes that a subset $M$ of features maintains the same dependency structure in both domains and hence has a consistent impact on the output $Y$ across domains.
The wenda method consists of three main components:
@@ 4 The Wenda Method
    Feature Models: wenda utilizes Bayesian models to estimate the dependency structure between the inputs in the source domain.
@@
    Confidence Scores: The estimated dependency structures are then evaluated in the target domain. This evaluation provides confidence scores for each feature, indicating how well it can be adapted for use in the target domain.
@@
    Final Adaptive Model: A final predictive model is trained using the source domain data. During training, the strength of regularization for each feature is adjusted based on its confidence score, which helps balance the relevance of all features for adaptation and their importance for the final prediction task on the target domain.
By applying stronger regularization to features with greater discrepancies between source and target domains, wenda allows for a trade-off between the suitability of features for domain adaptation and their predictive importance [6]. The results presented by Handl et al. demonstrate that this approach significantly minimizes the impact of distribution shifts, even in complex datasets from computational biology. Specifically, to showcase it performance, wenda was applied to age prediction from DNA methylation data. A key metric for assessing domain adaptation success, as highlighted by Handl et al., is the model’s accuracy on cerebellum samples, as this tissue is not included in the training set and is poorly predicted by a non-adaptive model. Their findings indicate that wenda significantly improves prediction accuracy for cerebellum samples compared to the standard non-adaptive model [6].
5 The Freda Method
Freda operates with the same underlying principles as wenda. Initially, the input dependency structure is collaboratively learned in a distributed setting by training feature models. This step is followed by the computation of confidence scores for each feature, and finally, the collaborative training of the adaptive model using the source domains. For simplicity, we explain our method with a single target domain; however, similar to Handl et al. [6], our method can be applied to multiple target domains, as demonstrated in Section 2.
A detailed overview of freda is presented in Figures 5 and 6. Figure 5 illustrates the first phase of our framework: feature model training with federated hyper-parameter optimization. This phase is repeated for each feature in the dataset. In this process, the source domain owners and the target domain owner split their data into two parts, excluding one feature at a time, where $X_{\neg f}$ represents the data after feature $f$ has been removed. The phase is completed once the closed-form solutions for all feature models are computed using secure aggregation and the FLAKE randomized encoding framework [13]. After obtaining predictions from the feature models, these predictions are shared with the target domain owner for the computation of confidence scores.
The subsequent phases of the framework are illustrated in Figure 6, which includes the federated optimal $\lambda$ prediction and the federated training of the final adaptive model. First, feature weights are computed from the confidence scores generated by the target domain owner in the previous step. Next, the target domain owner partitions the target data into two subsets, $X^{t}_{1}$ and $X^{t}_{2}$, where $X^{t}_{1}$ is used to predict the optimal $\lambda$ values required for training the final adaptive model.
For optimal $\lambda$ prediction, the source domain clients create a list of $j$ potential $\lambda$ values and collaboratively train separate models for each $\lambda$ on each domain in $X^{t}_{1}$. These models are then sent to the target domain owner, who selects the best-performing $\lambda$ for each domain in $X^{t}_{1}$. Using these selected $\lambda$ values and prior knowledge of domain similarities, the target domain owner predicts the optimal $\lambda$ values for the remaining domains in $X^{t}_{2}$. After this step, the target domain owner shares the predicted optimal $\lambda$ values with the source domain clients, enabling them to collaboratively train the final adaptive models. These models are then prepared for inference on the target domain data.
We now provide a detailed explanation of each of these steps.
Figure 5: Overview of the initial phases of the freda framework. The process begins with feature model training, which includes federated hyper-parameter optimization. For each feature, a closed-form solution is computed using the FLAKE randomized encoding framework. After obtaining the predictions from these feature models, they are shared with the target domain owner for the computation of confidence scores.
Figure 6: Subsequent phases of the freda framework. These include federated optimal lambda prediction and training of the final adaptive model. The target domain owner picks the optimal lambda values by splitting their data into two subsets, $X^{t}_{1}$ and $X^{t}_{2}$, and using the information from $X^{t}_{1}$ to infer the optimal lambda values for $X^{t}_{2}$. After sharing the optimal values with source domain clients, they collaboratively train the final models, which are then used for inference on the target domain data.
5.1 Feature Model Training
The performance of our method relies heavily on how well the input dependency structure between the source and the target domain data can be captured in a distributed setting without compromising data privacy. For this, we use Bayesian models to capture the dependency structure between the source and target domains, specifically employing Gaussian process regressors (GPR). For each feature $f$, we train a GPR model $g_{f}$ that predicts $f$ based on all other features using the source domain inputs $x_{1},\ldots,x_{n}$ as training data. These feature models estimate all conditional distributions $P_{S}(X_{f}\mid X_{\neg f})$.
Unlike other supervised regression models that predict a single value for a given input, GPR provides a full predictive distribution as the output [24], which we later use to compute confidence scores. Defining a GPR requires a covariance function, commonly referred to as a kernel, and the choice of kernel depends on the task and problem complexity. Given our focus on high-dimensional feature spaces, we use GPR models with a linear kernel and additive noise [25]. We define our linear kernel as follows:
$k(x,y)=\sum\limits_{i=1}^{D}\sigma_{p}^{2}x_{i}y_{i}$ (2)
Where $x$ and $y$ represent the elements in two sets for which the covariance function is to be computed, and $\sigma_{p}^{2}$ is the variance of the prior on the coefficients. Using this kernel, given training data $x=[x_{1},\ldots,x_{n}]$ with targets $y=[y_{1},\ldots,y_{n}]=[f(x_{1}),\ldots,f(x_{n})]$, with Gaussian noise in each point $\mathcal{N}_{0,\sigma_{n}^{2}}$ and new data points $x_{*}=[x^{*}_{1},\ldots,x^{*}_{m}]$ for which we want to predict $y_{*}=[f(x^{*}_{1}),\ldots,f(x^{*}_{m})]$, the closed-form solution of the GPR model can be written as follows [24]:
$\mathcal{N}(K_{*}K^{-1}y,K_{**}-K_{*}K^{-1}K_{*}^{T})$ (3)
where,
$K$ $=k(x_{i},x_{j})+\sigma_{n}^{2}*\mathbbm{1}_{n}$
$K_{*}$ $=k(x_{i},x_{j}^{*})$ (4)
$K_{**}$ $=k(x^{*}_{i},x^{*}_{j})$
This GPR model has two hyper-parameters that need to be optimized to achieve the best performance, the variance of the prior on the coefficients $\sigma_{p}^{2}$ in the covariance function shown in Eq. 2, and the variance of the additive noise $\sigma_{n}^{2}$ in the closed-form solution in Eq. 3. The optimal values for these two hyper-parameters are determined by maximizing the marginal likelihood for each feature. Similar to Eq. 3, given training data $x=[x_{1},\ldots,x_{n}]$ with their corresponding targets $y=[y_{1},\ldots,y_{n}]$, and the $K$ from Eq. 4 we maximize the following:
$\log\mathcal{L}(\sigma_{p}^{2},\sigma_{n}^{2})=-\frac{1}{2}y^{\top}K^{-1}y-\frac{1}{2}\log|K|-\frac{n}{2}\log(2\pi)$ (5)
Feature models are the most important component of freda, as the method relies on the input dependency structure captured by these models. Training feature models is straightforward when both target and source domains are accessible simultaneously. However, significant challenges arise when these datasets are distributed. The first challenge is that if the source domain is distributed across multiple entities, the optimization of hyper-parameters, shown in Eq. 5, cannot be performed across the entire source domain. The second, and more complex challenge, is that due to the distribution of the source and target domains, the closed-form solution of the GPR model (as shown in Eq. 3) cannot be computed directly. Since GPRs are non-parametric machine learning algorithms, to obtain predictions, one must compute the three matrices in Eq. 4, namely $K$, $K_{*}$, and $K_{**}$. In our setting, computing $K_{**}$ is straightforward and can be performed locally by the owner of the target domain, as it requires only target domain data. However, computing $K$ and $K_{*}$ is more challenging. Computing $K$ is challenging because, although it only requires source domain data, in our setting the source domain data is distributed across multiple entities and thus the entire matrix product of $K$ cannot be computed trivially and must be computed collaboratively among all the source domain owners while still preserving privacy. Whereas computing $K_{*}$ is challenging since it requires access to both $x$ (source domain) and $x_{*}$ (target domain). Freda addresses both of these challenges by employing secure aggregation and a special masking scheme for matrix product computation called FLAKE [13], enabling participants in the federated learning system to capture the dependency structure between domains with negligible difference from wenda, while ensuring complete data privacy.
5.1.1 Federated Hyper-Parameter Optimization
To optimize the hyper-parameters required for the GPR models in a distributed setting, specifically the variance of the prior on the coefficients $\sigma_{p}^{2}$ in the covariance function shown in Eq. 2 and the variance of the additive noise $\sigma_{n}^{2}$ in the closed-form solution in Eq. 3, we employ secure aggregation [14].
Secure aggregation is a broad term that encompasses various techniques aimed at computing a multiparty sum without revealing any participant’s update in plain-text, even to the aggregator [14]. While more sophisticated privacy-enhancing technologies such as Fully Homomorphic Encryption or Secure Multi-party Computation can be used for secure aggregation, we adopt a more practical approach called non-zero-sum masking.
Secure aggregation is a broad term that includes various techniques aimed at computing a multiparty sum without revealing any participant’s update in plain-text, even to the aggregator [14]. Although more complex privacy-enhancing technologies such as Fully Homomorphic Encryption or Secure Multi-party Computation can achieve secure aggregation, we employ a more practical approach called non-zero-sum masking.
Zero-sum masking, proposed by Bonawitz et al. [14], is a method for secure aggregation that enables participants to compute the sum of their inputs while preserving the privacy of each party’s data. Each party masks its data by adding a random value, ensuring that the masked values are indistinguishable from random noise. These random values are picked in such a way that, when the contributions from all participants are summed, the masking terms cancel out, revealing only the aggregated sum of the original data.
The process works as follows: for each party $i$, a random mask $r_{i}$ is generated and added to the data $d_{i}$, resulting in $d_{i}+r_{i}$. Each party then shares the masked data with the aggregator. At the same time, parties exchange the necessary random masks $r_{i}$ with one another so that, when summed, the masks $r_{i}$ cancel out across participants, leaving only the sum of the original data $d_{1}+d_{2}+\cdots+d_{n}$. The aggregator can thus compute the final sum without learning the individual contributions from any party.
This approach is efficient and avoids the computational overhead associated with more complex privacy-enhancing technologies making it well-suited for large-scale federated learning systems.
In freda, the optimization of the hyper-parameters $\sigma_{p}^{2}$ and $\sigma_{n}^{2}$ is carried out by the source domain data owners locally. Each source domain owner $m$ maximizes the marginal likelihood for each feature in their local data, following Eq. 5, to obtain ${\sigma_{p}^{2}}_{m}$ and ${\sigma_{n}^{2}}_{m}$. These locally optimized hyper-parameter pairs are then aggregated using zero-sum masking with the help of the aggregator. The goal is to estimate the global optimal hyper-parameters as if the marginal likelihood were maximized over the entire source domain data.
5.1.2 Federated GPR Computation
After the hyper-parameters for the feature models have been computed next step is to collaboratively compute the GPR models. In the GPR model prediction, the most challenging parts to compute are $K$ and $K_{*}$ as they require access to both all source domain data $X^{s}$ and target domain data $X^{t}$ to compute their matrix product in a naive plain-text approach. To handle the computation of $K$ and $K_{*}$ in a privacy preserving way, we utilize a secure and private matrix product computation framework, FLAKE [13]. It provides us with the product of the matrices from source domain and target domain without requiring access to the plaintext version of these matrices. FLAKE uses special masking matrices to hide the input matrices of the matrix product and reveals only the result of this matrix multiplication in the end. It protects the privacy of input matrices as well as their dimensionality. More specifically, both the parties with source domain and the party with target domain share a common seed to generate a common mask matrix $N\in\mathbb{R}^{d\times F}$ where $d$ is the dimensionality higher than the original space of the samples. After creating a common mask matrix $N$, each party $p$ in the computation locally computes a left inverse $L_{p}$ of $N$ where $L_{p}\in\mathbb{R}^{F\times d}$ such that $L_{p}\cdot N=I^{d}$. Using $L_{p}$ and $N$, each party $p$ computes its masked input matrix as follows:
$x_{p}^{{}^{\prime}}=x_{p}L_{p}(NN^{T})^{1/2}$
The parties send their masked data to the aggregator to allow the aggregator to compute the Gram matrix of the matrices of these parties, which is then used to compute the $K_{*}$. To calculate the Gram matrix, the aggregator performs the following computation of each pair of parties $p$ and $q$:
$\begin{split}x_{p}x_{q}&=x_{p}^{{}^{\prime}}(x_{q}^{{}^{\prime}})^{T}\\ &=(x_{p}L_{p}(NN^{T})^{1/2})(x_{q}L_{q}(NN^{T})^{1/2})^{T}\\ &=x_{p}L_{p}(NN^{T})^{1/2}((NN^{T})^{1/2})^{T}L_{q}^{T}x_{q}^{T}\\ &=x_{p}L_{p}(NN^{T})^{1/2}(NN^{T})^{1/2}L_{q}^{T}x_{q}^{T}\\ &=x_{p}L_{p}(NN^{T})L_{q}^{T}x_{q}^{T}\\ &=x_{p}(L_{p}N)(N^{T}L_{q}^{T})x_{q}^{T}\\ &=x_{p}II^{T}x_{q}^{T}\\ &=x_{p}x_{q}^{T}\end{split}$ (6)
Once the Gram matrix is computed, the aggregator can use the entries of this matrix to compute $K$ and $K_{*}$ as follows:
$\begin{split}K&=k(x^{p},x^{q})+\sigma_{n}^{2}*\mathbbm{1}_{n}\\ &=\sigma^{p}\sum_{i}^{D}x^{p}_{i}y^{q}_{i}+\sigma_{n}^{2}*\mathbbm{1}_{n}\\ &=\sigma^{p}G_{pq}+\sigma_{n}^{2}*\mathbbm{1}_{n}\\ \end{split}$
and
$\begin{split}K_{*}&=k(x^{p},x^{q})\\ &=\sigma^{p}\sum_{i}^{D}x^{p}_{i}y^{q}_{i}\\ &=\sigma^{p}G_{pq}\\ \end{split}$
This process allows the parties to collaboratively compute $K$ and $K_{*}$ without revealing their data to the aggregator. Computing the missing components, the aggregator can now compute the closed-form solution of the optimization function given in Equation 3 using $K$ and $K_{*}$ along with $K_{**}$ and the label vectors, leading directly to the predictions of the feature models required for the computation the confidence scores.
5.2 Confidence Score Computation

### v2

@@ 3 Discussion
In this study, we introduce freda, the first method to enable privacy-preserving, federated unsupervised domain adaptation for regression tasks in high-dimensional, small-scale biological datasets.
Existing FDA methods predominantly rely on deep learning models, which require large datasets and often struggle with high-dimensional, small-sample-size biological data. Furthermore, these methods primarily focus on classification tasks, leaving regression-based FDA approaches largely underexplored. In contrast, freda introduces a novel combination of randomized encoding and secure aggregation to enable the first privacy-preserving, federated training of Gaussian Processes. By leveraging Gaussian Processes, freda effectively models complex feature relationships in high-dimensional small-scale datasets while preserving complete data privacy, making it particularly well-suited for biomedical applications.
We evaluate freda on a challenging benchmark task of age prediction from DNA methylation data. Our experimental results demonstrate that freda performs comparably to centralized state-of-the-art domain adaptation method for small-scale and high-dimensional datasets while significantly outperforming the non-adaptive centralized baseline, even when source domain data is distributed across multiple parties. Notably, for cerebellum samples, which exhibit a challenging distribution shift, freda effectively models complex feature relationships and adapts to the shift without requiring raw data sharing. This highlights its potential for real-world applications where data heterogeneity and privacy constraints pose significant challenges.
A key contribution of our work is showing that privacy-preserving federated training of Gaussian Processes is not only feasible but also highly effective for distributed domain adaptation in biological datasets. Traditional Gaussian Process models, while well-suited for feature modeling, have never been scaled to distributed settings due to their reliance on pairwise data computations. Freda is the first method to overcome this limitation by enabling secure and collaborative training, ensuring that institutions can benefit from improved feature modeling without compromising data privacy.
4 Related Work
4.1 Domain Adaptation
Domain adaptation, particularly in the unsupervised setting, has been extensively studied in recent years [25, 26, 27]. The goal of domain adaptation is to transfer knowledge from a source domain to a target domain, where labeled data in the target domain is scarce or unavailable [28]. While numerous methods exist, the majority of these focus on image classification problems, where the source data is typically abundant and the number of features is relatively small.
Notable unsupervised domain adaptation techniques often aim to align feature representations across domains. One common approach involves adversarial training to minimize the domain discrepancy. For instance, the Domain-Adversarial Neural Network proposed by Ganin et al. [29] uses a gradient reversal layer to align source and target feature distributions via adversarial learning, significantly improving classification performance on tasks like digit recognition. Another popular method is Maximum Mean Discrepancy based alignment, proposed by Long et al.’s Deep Adaptation Networks, which align the distributions of deep features between domains using kernel-based metrics [30]. Additionally, self-supervised learning approaches such as CDAN [31] integrate task-specific predictions into domain alignment to capture conditional dependencies. These methods are widely adopted in benchmarks like Office-31, VisDA, or digit datasets, showcasing high performance in handling distribution shift in image classification.
4.2 Domain Adaptation for Regression
Unlike classification problems, domain adaptation for regression has received less attention over the years, primarily due to the challenges in aligning continuous output spaces. The recently proposed Distribution-Informed Neural Network (DINO) [32] framework introduces a unified approach to tackle this issue by leveraging distribution-aware neural networks. Empirical results demonstrate that DINO achieves state-of-the-art performance across domain adaptation regression benchmarks, highlighting its effectiveness in mitigating distribution shifts in regression tasks [32]. Still, DINO has been tested on image datasets with high sample sizes and low feature counts, while effective in such scenarios, these datasets are not representative of the high-dimensional, low-sample-size challenges faced in biological datasets.
A key state-of-the-art method for unsupervised domain adaptation in high-dimensional, small-sample datasets is wenda [20]. Unlike deep learning-based approaches that require large datasets, wenda is specifically designed for tabular biological data and operates under a relaxed covariate shift assumption, allowing for domain adaptation even when certain features behave differently across domains. It achieves this by estimating feature dependencies in the source domain, assessing their consistency in the target domain, and applying adaptive regularization to account for domain shifts. In their study, Handl et al. demonstrated that wenda significantly improves adaptation performance in DNA methylation-based age prediction, particularly for cerebellum samples, which exhibit substantial distributional shifts. Due to its effectiveness in this domain, we use wenda as a baseline method to compare with freda.
4.3 Federated Domain Adaptation
Federated domain adaptation extends the principles of domain adaptation to the FL setting, where data is located on decentralized devices or institutions, and privacy constraints prevent data sharing. In this context, methods must tackle both domain shifts, the heterogeneity inherent in FL environments, and the privacy concerns. Multiple federated domain adaptation techniques have been proposed in the recent years. For instance, PartialFed [33] introduces a strategy where only a subset of global model parameters is loaded on local clients, allowing for dynamic mixing of global and client-specific parameters. This approach enhances performance on cross-domain classification and detection tasks, demonstrating significant improvements on image datasets such as Office-Home and UODB [33].
Similarly, FedGP [34] proposes a gradient-projection-based aggregation method to enhance the performance of the model on a target client with limited data. By filtering out noisy source gradients and optimally combining target and source gradients, FedGP achieves notable performance improvements under both domain shift and data scarcity scenarios. Experiments on real-world image datasets verify its robustness and efficacy in FDA settings [34].
Still, these studies primarily focus on image datasets with abundant samples and relatively low-dimensional features, which may not translate effectively to domains such as computational biology, where data is high-dimensional and samples are limited.
5 The Freda Method
We consider the following distributed setting: there are $N$ source domain clients, each with a local labeled dataset $X^{s_{i}}=\{(x^{s_{i}}_{j},y^{s_{i}}_{j})\}_{j=1}^{n_{i}}$ and a sample size of $n_{i}$. The entire source domain data, distributed across the $N$ clients, is denoted as $X^{S}=\bigcup_{i=1}^{N}X^{s_{i}}$. Similarly, there is a target client with a dataset $X^{T}=\{x^{t}_{m}\}_{m=1}^{n_{t}}$ containing $n_{t}$ samples for the same prediction task, but without any available labels. In both the source and target domain datasets, the samples $\{x^{s_{i}}_{j}\}$ and $\{x^{t}_{m}\}$ are $\mathcal{P}$-dimensional vectors, where $\mathcal{P}\in\mathbb{N}$, and the labels $\{y^{s_{i}}_{j}\}$ are scalars.
The goal of our framework is to leverage the inputs and labels from the source domains, along with the inputs from the target domain, to develop a model that achieves high performance on the target domain but without requiring explicit data sharing between parties.
Freda operates with the same underlying principles as wenda. The process begins with collaboratively learning the input dependency structure in a distributed setting by training feature models. Next, confidence scores are computed from the predictions of these feature models, which are then transformed into feature weights. Subsequently, optimal lambda prediction is performed by federatively training multiple models on the source domain using varying regularization parameters. This step enables the target domain owner to determine the optimal regularization parameter. Finally, the method concludes with the federated training of the final adaptive model on the source domains.
For simplicity, we describe our method using a single domain in the target data; however, similar to the approach of Handl et al. [20], Freda can be extended to multiple target domains inside the target data, as demonstrated in Section 2.
5.1 Feature Model Training
The performance of our framework relies on capturing the input dependency structure between the source and target domain data in a distributed setting while maintaining data privacy. To achieve this, we leverage Bayesian models, specifically Gaussian Process Regressors, to model the dependency structure. For each feature $f$, we train a GPR model $g_{f}$ that predicts $f$ based on all other features, using the source domain inputs from all source clients as training data.
The training data for a given feature $f$ includes $X^{S}_{\neg f}$, the entire source domain data with the column corresponding to feature $f$ removed, as inputs, and $X^{S}_{f}$, the corresponding feature vector, as labels. For new data points $X^{T}_{\neg f}=[x^{t}_{1,\neg f},\ldots,x^{t}_{n_{t},\neg f}]$, the target domain data with the column for feature $f$ removed, the goal is to predict the corresponding feature vector $X^{t}_{f}=[x^{t}_{1,f},\ldots,x^{t}_{n_{t},f}]$.
For a specific source domain $i$, the vector containing feature $f$ is denoted $X^{s_{i}}_{f}=[x^{s_{i}}_{1,f},\ldots,x^{s_{i}}_{n_{i},f}]$, while the $(n_{i}\times(\mathcal{P}-1))$-matrix of remaining features is $X^{s_{i}}_{\neg f}=[x^{s_{i}}_{1,\neg f},\ldots,x^{s_{i}}_{n_{i},\neg f}]$. Thus, for a given feature $f$, the GPR model $g_{f}$ provides a closed-form predictive distribution:
$\mathcal{N}(K_{*}K^{-1}y,K_{**}-K_{*}K^{-1}K_{*}^{\top})$ (2)
where:
$K$ $=k(X^{S}_{\neg f},X^{S}_{\neg f})+\sigma_{n}^{2}\mathbbm{1}_{n_{S}}$
$K_{*}$ $=k(X^{S}_{\neg f},X^{T}_{\neg f})$ (3)
$K_{**}$ $=k(X^{T}_{\neg f},X^{T}_{\neg f})$
$y$ $=X^{S}_{f}$
Here, $k(.,.)$ computes the linear kernel with the variance of the prior on the coefficients $\sigma_{p}^{2}$ between the given input matrices [35]. Additionally, $n_{S}$ denotes the total number of samples in the entire source domain dataset $X^{S}$. Unlike traditional supervised regression models that predict a single value for a given input, GPRs provide a full predictive distribution as output [36], which we later use to compute feature weights. This GPR model involves two hyper-parameters that must be optimized for the best performance: the variance of the kernel, $\sigma_{p}^{2}$, and the variance of the additive Gaussian noise, $\sigma_{n}^{2}$, from the closed-form solution in Eq. 2. The optimal values of these hyper-parameters are determined by maximizing the marginal likelihood for each feature. For a specific source client $i$, and the covariance matrix $K=k(X^{s_{i}}_{\neg f},X^{s_{i}}_{\neg f})+\sigma_{n}^{2}\mathbbm{1}_{n_{i}}$ from Eq. 3, source client $i$ maximizes:
$\log\mathcal{L}(X^{s_{i}}_{f}|X^{s_{i}}_{\neg f})=-\frac{1}{2}(X^{s_{i}}_{f})^{\top}K^{-1}X^{s_{i}}_{f}-\frac{1}{2}\log|K|-\frac{n_{i}}{2}\log(2\pi)$ (4)
Training feature models is straightforward when both target and source domains are accessible simultaneously. However, significant challenges arise when these datasets are distributed.
The first challenge is that, if the source domain is distributed across multiple entities, the optimization of hyper-parameters shown in Eq. 4 cannot be performed across the entire source domain. The second, and more complex, challenge is that due to the distribution of the source and target domains, the closed-form solution of the GPR model (as shown in Eq. 2) cannot be computed directly. Since GPRs are non-parametric machine learning algorithms, obtaining predictions requires computing the three matrices in Eq. 3, namely $K$, $K_{*}$, and $K_{**}$.
In our setting, computing $K_{**}$ is straightforward and can be performed locally by the owner of the target domain, as it requires only the target domain data. However, computing $K$ and $K_{*}$ as well as the predictive mean of the feature model $K_{*}K^{-1}y$ presents significant challenges. Computing $K$ is challenging because, although it only requires source domain data, the data is distributed across multiple entities, and the entire matrix product of $K$ cannot be computed trivially. Instead, it must be computed collaboratively among all source domain owners while preserving privacy. Similarly, computing $K_{*}$ is challenging because it requires access to both source domain data and target domain data. Lastly, computing $K_{*}K^{-1}y$ is not straightforward because the feature column for the entire source domain data, $y=X^{S}_{f}$, is distributed between $N$ source clients. Moreover, explicitly sharing $y$ would compromise privacy since the aggregator could reconstruct the data of all source clients after the feature model training phase is completed, as each feature is modeled independently.
Freda addresses all of these challenges by employing secure aggregation and a special masking scheme for matrix product computation called FLAKE [37].
5.1.1 Federated Hyper-Parameter Optimization
To optimize the hyper-parameters required for the GPR models in a distributed setting, specifically the variance of the prior on the coefficients $\sigma_{p}^{2}$ in the kernel function and the variance of the additive noise $\sigma_{n}^{2}$ in the closed-form solution (Eq. 2), we employ secure aggregation [38].
Secure aggregation refers to a range of techniques designed to compute a multiparty sum without exposing any individual participant’s data in plain text, even to the aggregator [38]. While advanced privacy-enhancing technologies like Fully Homomorphic Encryption (FHE) or Secure Multi-party Computation (MPC) can achieve secure aggregation, Freda adopts a more practical approach using a method known as zero-sum masking.
Zero-sum masking, as proposed by Bonawitz et al. [38], allows participants to securely compute the sum of their inputs while ensuring the privacy of each party’s data. In this approach, each participant masks its data with a random value, making the masked values indistinguishable from random noise. The random values are chosen such that, when the contributions from all participants are summed, the masking terms cancel out, revealing only the aggregated sum of the original data.
The process is as follows: for each party $i$, a random mask $r_{i}$ is generated and added to the data $d_{i}$, resulting in $d_{i}+r_{i}$. Each party shares its masked data with the aggregator while simultaneously exchanging the necessary random masks $r_{i}$ with other parties. When the aggregator computes the sum of all masked values, the random masks cancel out across participants, revealing only the aggregated sum $d_{1}+d_{2}+\cdots+d_{N}$. This method ensures data privacy without requiring complex cryptographic operations, making it efficient and suitable for large-scale federated learning systems.
In Freda, the optimization of the hyper-parameters $\sigma_{p}^{2}$ and $\sigma_{n}^{2}$ is performed locally by each source client. Each source client $i$ maximizes the marginal likelihood for every feature in its local dataset $\{x^{s_{i}}_{j}\}_{j=1}^{n_{i}}$, following Eq. 4, to obtain $\sigma_{p,i}^{2}$ and $\sigma_{n,i}^{2}$. These locally optimized hyper-parameter pairs are then aggregated using zero-sum masking, enabling the aggregator to compute global averages for $\sigma_{p}^{2}$ and $\sigma_{n}^{2}$ as if the marginal likelihood had been maximized over the entire source domain data.
5.1.2 Federated GPR Computation
In the GPR model prediction process, the most challenging components to compute are $K$, $K_{*}$, and their product with the global feature vector $y=X^{S}_{f}$, as these require access to both the source domain data and the target domain data to calculate the matrix product. A naive plaintext approach would compromise privacy. To overcome this, we utilize FLAKE [37], a framework for secure and private matrix product computation, which allows us to calculate the product of matrices from the source and target domains without disclosing their plaintext values.
Privacy-Preserving Masking Process
We use special masking matrices to hide the input matrices of the matrix product and reveal only the result of this multiplication to the aggregator. This protects the privacy of the input matrices as well as their dimensionality. Specifically, both the source clients and the target client share a common seed to generate a shared mask matrix $M\in\mathbb{R}^{d\times\mathcal{P}}$, where $d$ is a higher-dimensional space than the original feature space. Each client $p$ locally computes a left inverse $L_{p}\in\mathbb{R}^{\mathcal{P}\times d}$ for $M$, such that $L_{p}M=I_{d}$, and applies the mask to its data:
$x_{p}^{\prime}=x_{p}L_{p}(MM^{\top})^{1/2},$
where $x_{p}$ represents the local dataset of client $p$ ($x_{p}=X^{s_{i}}$ for a source client or $x_{p}=X^{T}$ for the target client). The masked data $x_{p}^{\prime}$ is sent to the aggregator.
Secure Gram Matrix Computation
The aggregator computes the Gram matrix for all clients using the masked data. For a pair of clients $p$ and $q$, the Gram matrix is computed as:
$\begin{split}x_{p}x_{q}^{\top}&=x_{p}^{{}^{\prime}}(x_{q}^{{}^{\prime}})^{\top}\\ &=(x_{p}L_{p}(MM^{\top})^{1/2})(x_{q}L_{q}(MM^{\top})^{1/2})^{\top}\\ &=x_{p}L_{p}(MM^{\top})^{1/2}((MM^{\top})^{1/2})^{\top}L_{q}^{\top}x_{q}^{\top}\\ &=x_{p}(L_{p}M)(M^{\top}L_{q}^{\top})x_{q}^{\top}\\ &=x_{p}x_{q}^{\top},\end{split}$ (5)
where the masking terms cancel out, revealing only the product $x_{p}x_{q}^{\top}$ without exposing the individual data of the clients.
Using the computed Gram matrix $G_{pq}$, the aggregator calculates $K$ and $K_{*}$ as follows:
$K$ $=\sigma_{p}^{2}G_{pq}+\sigma_{n}^{2}\mathbbm{1}_{n_{S}},$
$K_{*}$ $=\sigma_{p}^{2}G_{pq}.$
Computing the Predicted Mean
Computing the predicted mean $K_{*}K^{-1}y$ is challenging as the aggregator does not have access to the label vector $y$ (feature column for the modeled feature $g_{f}$), which remains distributed across source clients. Explicitly sharing $y$ would also compromise privacy since the aggregator could reconstruct the data of all source clients after the feature model training phase is completed, as each feature is modeled independently. To address this, Freda employs randomized encoding.
@@ 5.1.2 Federated GPR Computation
    Aggregator Step: The aggregator computes the masked intermediate result $CK_{*}K^{-1}$, where $C\in\mathbb{R}^{n_{t}\times n_{t}}$ is a random mask matrix, and sends $C^{-1}$ to the target client.
@@
    Source Client Collaboration: The masked matrix $CK_{*}K^{-1}$ is split into sub-matrices corresponding to each source client’s data represented as $(CK_{*}K^{-1})^{s_{i}}$. Each source client $i$ computes:
    $(CK_{*}K^{-1})^{s_{i}}X^{s_{i}}_{f},$
    This computation results in a $n_{t}\times 1$ vector, which each source client sends to the target client.
@@
    Target Client Aggregation: The target client aggregates the received vectors and removes the mask using $C^{-1}$:
    $K_{*}K^{-1}X^{S}_{f}=C^{-1}\sum_{i=1}^{N}(CK_{*}K^{-1})^{s_{i}}X^{s_{i}}_{f}.$
This process enables the target client to compute the predicted mean while ensuring that the privacy of individual source clients’ data is preserved. Additionally, it protects against source clients inferring any information from the sub-matrices of the intermediate matrix product $K_{*}K^{-1}$.
Computing the Predicted Variance
The predicted variance $K_{**}-K_{*}K^{-1}K_{*}^{\top}$ is computed using the Gram matrix $G_{pq}$ and sent to the target client by the aggregator. With both the predicted mean and variance, the target client obtains the complete closed-form solution of the GPR model.
5.2 Feature Weight Computation